A cliff is 20 m above the ground. A boy rolls a ball off the edge with an unknown initial velocity and it strikes the ground a distance of 3 m horizontally away. Calculate (a) the ball’s initial velocity, (b) its time of flight, and (c) the magnitude and direction of its velocity when it strikes the ground


1
Expert's answer
2021-10-01T17:52:35-0400

a)

"t=\sqrt{\frac{2h}g}=2~s,"

b)

"v_x=\frac lt=1.5~\frac ms,"

c)

"v_y=\sqrt{2gh}=20~\frac ms,"

"v=\sqrt{v_x^2+v_y^2}=20.056~\frac ms,"

"\alpha=\arctan\frac{v_y}{v_x}=85.7°."